If it has been determined that the probability of an earthquake occurring on a certain day in a certain area is 0.03125, what are the odds AGAINST an earthquake?
Answer by priya(51)

In Probability,
P(A) + P(A') = 1
Where P(A) is the probability of a thing to happen and P(A') is the proability of that thing does not happen.

so the probability of earthquake not happening is given by,
0.03125 +P(A') =1
P(A') = 1 - 0.01325
=0.96875 IS THE ANSWER.
VISIT MY SITE (CLICK ME) TO LEARN MATHS ONLINE..
PRIYA